# Language Models are Crossword Solvers

### v1

@@ 1 Introduction
There is another aspect to solving crossword puzzles, which is constraint satisfaction (see Figure 1). Not only does the generated output have to adhere to the length limitations, it also has to incorporate character level constraints from previously generated answers. Successful crossword solvers generate answers based on clues, length information, character information from earlier answers, etc. and have the ability to backtrack when faced with inconsistent solutions in a targeted manner. Thus, even when the crossword clues are simpler, as in straight crosswords, the problem still poses a significant challenge to SoTA LLMs.
Since solving crosswords requires proficiency in several desirable areas, our objective with this paper is to probe LLMs’ abilities at this multi-faceted task. To that end, we first perform extensive analysis to understand how well LLMs can solve crossword clues based on the provided contextual information and constraints. We show that current-generation SoTA LLMs demonstrate massively improved performance when compared to previous SoTA baselines, often outperforming them by a factor of 2 or 3, without any fine-tuning.
@@
Armed with the knowledge gathered from our analysis, we tackle the full problem - i.e. not only answering clues, but filling in the grid considering all the imposed constraints. We device an algorithm called SweepClip which can solve straight crossword puzzles with an accuracy of 93.1% with the aid of an LLM. This algorithm exploits the constraints imposed by previously generated answers to improve future answers, and achieves results comparable to rule-based systems. This algorithm also improves the clue deciphering capability of LLMs significantly.
@@ 2 Background
Traditional approaches to solving crossword puzzles typically involve two parts - a candidate answer proposal system and a grid-filling algorithm (Angelini et al., 2005; Ginsberg, 2011). One such system, Proverb, Littman et al. (2002) achieves a letter accuracy of $98.1\%$ on crosswords from the New York Times. The candidate answer proposal systems rely on large datasets of crossword clues and answers (Shazeer et al., 1999) and grid-filling relies on algorithms for the constraint satisfaction problem (Angelini et al., 2005; Zugarini et al., 2023; Shazeer et al., 1999).
There have been attempts to solve cryptic crossword clues with traditional algorithms, employing a large dataset of clues and a curated context free grammar (CFG) based parser (Deits, 2015). However, this approach was not found to be performant ($7\%$ accuracy according to Sadallah et al. (2024)).
Using LLMs to solve cryptic crossword clues has received attention in some recent papers. Efrat et al. (2021) curated a large dataset of cryptic crossword clues sourced from two UK dailies - The Times and The Telegraph and fine-tuned a T5-Large (Raffel et al., 2020) model to establish baseline performances. They proposed a training split for their dataset where the answers in the training and test sets are mutually exclusive so the model cannot rely on memorization. Following this, we have the work by Rozner et al. (2021), where they curated a dataset of cryptic crossword clues from The Guardian, and fine-tuned a T5-Large model with a curriculum learning approach, demonstrating performance improvements. Rozner et al. (2021) had some salient criticisms of the work by Efrat et al. (2021), and suggested that a disjoint train-test split is not a strong enough condition to ensure that the models learn to solve cryptic crosswords, since the models demonstrate “…robustness to plural and other inflections”. Rozner et al. (2021) proposed a split where words with similar roots are grouped together, and demonstrated that this stringent criteria leads to reduced performance.
@@
All three recent works on cryptic crosswords conclude that there is a massive gap in performance between LLMs and human experts. Also, these recent works focusing on solving cryptic crosswords with LLMs do not attempt to solve the full problem, i.e. they do not attempt to incoporate grid information and the constraints imposed by them, and rather elect to pose the problem of solving crosswords in the form of a question answering (QA) task. In this paper, we undertake the challenge of solving a full crossword puzzle by developing an algorithm that incorporates grid information based on the answers given by an LLM. We find that with the application of our algorithm, the base performance of LLMs can be boosted significantly, and that the gap between the performance of LLMs and human experts is much smaller than previously anticipated. With future work addressing some key limitations, LLMs will compare favourably to human experts in this task.
@@ 3.1 Datasets
Most of our analysis is performed on three crossword puzzle datasets. The first one covering straight or American style crosswords was curated by us from the very popular and long running puzzle section of the New York Times. The other two, namely Cryptonite by Efrat et al. (2021) and word-init-disjoint by Rozner et al. (2021) cover cryptic crossword puzzles. As discussed in Section 2, the two cryptic crossword datasets have methodological differences in how they partition training-test splits, however, this is not a very pertinent consideration for us, since we do not perform any training. A bulk of the results were reported on the New York Times dataset we curated (abbreviated as NYT) for straight crosswords and word-init-disjoint (abbreviated as Init) for cryptic crosswords since it was found to be more challenging than Cryptonite for LLMs.
We also curate smaller additional datasets after the knowledge cut-off date of LLMs from The Guardian, and Lovatts Puzzles for analyzing generalizability and data contamination in LLMs. We have dedicated Appendix B to outline in detail all the protocols involved in curating and handling the data.
@@ 3.2 Answering Crossword Clues
The first step in solving a crossword puzzle is deciphering its clues, so following in the scheme of Rozner et al. (2021); Sadallah et al. (2024) and Efrat et al. (2021) we first explore this as a QA task, where the clue and its length is provided to the LLM as a query, with the expectation that it will produce the corresponding answer.
@@
We find that the performance difference between 5-shot and 10-shot prompted answers is not significant (see Figure 2), which is consistent with the findings by Sadallah et al. (2024) [^3] We elected not to perform experiments with 10-shot prompts on Claude and GPT-4-Turbo due to budget constraints.. We also note that there is an appreciable difference (up to 5%) in performance of models across the board between Cryptonite and Init, with LLMs showing diminished performance on the Init dataset.
@@ 4 SweepClip - Our proposed algorithm to solve crosswords with LLMs
Our proposed algorithm (outlined in Algorithm 1 and detailed in Appendix A), first generates a set of candidate answers for all clues provided with the crossword (sweep) and uses a graph based criteria to eliminate answers that do not fit well to the grid (clip). Following, this we use the constraints generated from the previous step answers, and generate more candidate answers, and prune the bad fitting candidates. We iteratively apply this strategy, until either - (i) the entire crossword is filled, (ii) we exceed a preset number of iterations, or (iii) we run out of LLM computational budget.
For pruning we use the largest connected component from the answers generated so far to ensure that the generated answers are somewhat coherent amongst themselves. This is over-restrictive, as it is possible that isolated answers are correct, however we find that this strict pruning strategy works better to eliminate bad answers early rather than using the constraints imposed by potentially bad answers to generate further bad answers.
We would like to highlight that our algorithm is sub-optimal, and does not provide correctness guarantees beyond elimination of conflicting answers, however, it was designed from the ground up to minimize cost incurred from calls to the LLM by avoiding branching based on LLM answers. We could not explore standard graph search algorithms or back-tracking to thorougly explore the consequences of generated answers, as that would necessitate a much larger computational budget. Future works attempting this with more resources would definitely see further performance gains by either running our algorithm with higher budgets or more thoroughly exploring the consequences of many possible candidate answers before selecting a final answer.
@@ 5.2 Solving Crossword Puzzles
For this task we employed our algorithm SweepClip on a set of 100 randomly sampled Monday New York Times crossword puzzles. We used two LLMs for this GPT-4-Turbo with max_depth of 30 and a budget of 0.5 USD per crossword, and LLaMA 3 70B with max_depth of 35 and a budget of 600 LLM calls.
@@
We find (see Table 4) that our algorithm with GPT-4-Turbo solves 48% of crosswords without any errors, and 55% of crosswords with at most 1 character wrong. The average character level accuracy in crossword solving is 93.1% ($\pm$ 14.1%). Our algorithm improves the clue-wise answer accuracy888clue level accuracy is different from character level, as it is possible to have a filled in crossword without deciphering all clues. (exact match) to 89.6% ($\pm$ 16.9%) from the base accuracy (without the algorithm) which is 43.5% ($\pm$ 23.5%).
@@
Our algorithm implicitly performs self-consistency checks to improve candidate answers. For example the generated clue at the first sweep may be discarded and regenerated with a different prompt and then accepted if it does not produce conflicts, and is coherent with a larger number of other answers. Thus with the application of our algorithm, we have successfully exploited the constraint information, to boost the performance of LLMs far beyond what would have been possible with straightforward QA like clue deciphering. To the best of our knowledge this is the first algorithm that demonstrates succesfully solving crosswords with the aid of LLMs.
@@ 5.3 Data Contamination and Generalizability
Having seen remarkable performance improvements in the current generation of SoTA LLMs, we set out to investigate the reason behind this success. To rule out data-contamination, we curated additional cryptic crossword datasets comprising entirely of puzzles published after May 20, 2024, which is after the knowledge cutoff date of all LLMs examined (see Table 5).  [^9] More details on data curation is in Appendix B.
@@
Note, the Init dataset is also sourced from the The Guardian, thus the results reported in Table 5 should be consistent. We see no appreciable difference in performance on the post-cutoff dataset, leading us to conclude that these LLMs can generalize beyond potential contamination in their training set. We also curate a dataset from Lovatts Puzzles  [^10] https://lovattspuzzles.com (post-cutoff), and observe even better performance.
To ascertain if the models can reason about unseen cryptic crossword clues to arrive at correct responses, we elect to perform human evaluation of model responses. We employ a 3-shot Chain-of-thought prompt to elicit a reasoned response to crossword clues from GPT-4-Turbo, which are then analyzed by our team for soundness vis-à-vis factual and logical errors. We choose 100 samples from the post-cutoff Lovatts set and rely on consensus of all evaluators to ascertain if a response contains factual or logical errors (e.g. wrong counting, wrong anagram, drawing a conclusion that does not follow from the premise).
@@ 6 Conclusions and Discussions
Solving crosswords require proficiency in a multitude of desirable skills, and cryptic crosswords was previously thought to be firmly in the domain of human experts. Our findings challenge this notion. We have demonstrated that the current generation of SoTA LLMs show significantly improved aptitude at solving straight and cryptic crossword puzzle clues, and can exploit constraints provided by partially solved crosswords to boost this performance further. We also found that this emergent ability generalizes well to the post knowledge cut-off regime and is accompanied with the capacity to produce reasoned explanations.
@@
This calls into question previously reported results which claim that LLMs massively lag behind human experts at solving cryptic crosswords. We are firmly of the opinion that with the improvements brought forth, the next generation of LLMs coupled with a a well-designed search algorithm will solve this “unsolvable” task. Incorporating length constraints effectively still remains a challenge, and further enhancements in numerical reasoning abilities in LLMs is a promising avenue to pursue.
@@
Limitation 2 - Algorithm We devised a new algorithm, when perhaps standard graph traversal algorithms would have performed better. Our algorithm potentially discards correct answers and does not fully explore the consequences of each possible generated answer. This is again a computational cost saving technique. Ideally, we would generate one candidate answer, followed by all its neighbors, and whenever there is a conflict, branch out and explore all options to see which is a better fit. We elect not to do this, because it involves a potentially exponential number of LLM calls. Future studies with much larger computational budgets will definitely have more success employing such a strategy.
@@ C.1 General Overview
Phi 3 mini Instruct 3.8B 4K October 2023
Mistral v0.2 Instruct 7B 32K December 2023
LLaMA 3 Instruct 8B 8K March 2023
Mixtral v0.1 8x7B 32K December 2023
LLaMA 2 70B 4K September 2022
LLaMA 3 70B 8K December 2023
Claude 3 sonnet 20140229 ? 200K March 2024
GPT-3.5-Turbo (0125) ? 16K September 2021
@@
We used temperature=0.5 for all models and experiments. For few-shot prompt responses max-tokens were set to 10, and for Chain-of-thought responses max-tokens was set to 1000. top_p, top_k or any other parameters were not changed. The Jinja templates used to format the text prompts are provided alongside the paper, and are borrowed from the model-cards of the respective models. Claude and GPT models were accessed through their respective APIs and all other models were run locally on one server consisting of 2x80GB Nvidia A100 GPUs.

### v2

@@ 1 Introduction
There is another aspect to solving crossword puzzles, which is constraint satisfaction (see Figure 1). Not only does the generated output have to adhere to the length limitations, it also has to incorporate character level constraints from previously generated answers. Successful crossword solvers generate answers based on clues, length information, character information from earlier answers, etc. and have the ability to backtrack when faced with inconsistent solutions in a targeted manner. Thus, even when the crossword clues are simpler, as in straight crosswords, the problem still poses a significant challenge to SoTA LLMs. Since solving crosswords requires proficiency in several desirable areas, our objective with this paper is to probe LLMs’ abilities at this multi-faceted task.
To that end, we first perform extensive analysis to understand how well LLMs can solve crossword clues based on the provided contextual information and constraints. We show that current-generation SoTA LLMs demonstrate massively improved performance when compared to previous SoTA baselines, often outperforming them by a factor of 2 or 3, without any fine-tuning.
@@
Armed with the knowledge gathered from our analysis, we tackle the full problem - i.e. not only answering clues, but filling in the grid considering all the imposed constraints. We device an algorithm called SweepClip which can solve straight crossword puzzles with an accuracy of 93.1% with the aid of an LLM. This algorithm exploits the constraints imposed by previously generated answers to improve future answers, and achieves results comparable to rule-based systems. This algorithm also improves the clue deciphering capability of LLMs, more than doubling the initial performance.
@@ 2 Background
Traditional approaches to solving crossword puzzles typically involve two parts - a candidate answer proposal system and a grid-filling algorithm (Angelini et al., 2005; Ginsberg, 2011). One such system, Proverb, Littman et al. (2002) achieves a letter accuracy of $98.1\%$ on crosswords from the New York Times (NYT). The candidate answer proposal systems rely on large datasets of crossword clues and answers (Shazeer et al., 1999) and grid-filling relies on algorithms for the constraint satisfaction problem (Angelini et al., 2005; Zugarini et al., 2023; Shazeer et al., 1999; Kulshreshtha et al., 2022). Kulshreshtha et al. (2022) who curated a dataset of NYT crossword puzzles and established benchmarks with language models have highlighted this task as “… a new high bar for AI systems”.
Attempts to solve cryptic crossword clues with traditional algorithms, employing a large dataset of clues and a curated context free grammar (CFG) based parser (Deits, 2015), were not found to be performant ($7\%$ accuracy according to Sadallah et al. (2024)). Using LLMs to solve cryptic crossword clues has received attention in some recent papers. Efrat et al. (2021) curated a large dataset of cryptic crossword clues sourced from two UK dailies - The Times and The Telegraph and fine-tuned a T5-Large (Raffel et al., 2020) model to establish baseline performances. They proposed a training split for their dataset where the answers in the training and test sets are mutually exclusive so the model cannot rely on memorization. Following this, we have the work by Rozner et al. (2021), where they curated a dataset of cryptic crossword clues from The Guardian, and fine-tuned a T5-Large model with a curriculum learning approach, demonstrating performance improvements. Rozner et al. (2021) had some salient criticisms of the work by Efrat et al. (2021), and suggested that a disjoint train-test split is not a strong enough condition to ensure that the models learn to solve cryptic crosswords, since the models demonstrate “…robustness to plural and other inflections”. Rozner et al. (2021) proposed a split where words with similar roots are grouped together, and demonstrated that this stringent criteria leads to reduced performance.
@@
All three recent works on cryptic crosswords conclude that there is a massive gap in performance between LLMs and human experts. Also, these recent works focusing on solving cryptic crosswords with LLMs do not attempt to solve the full problem, i.e. they do not attempt to incoporate grid information and the constraints imposed by them, and rather elect to pose the problem of solving crosswords in the form of a question answering (QA) task. Kulshreshtha et al. (2022) attempted to solve NYT crossword puzzles with language models and a Satisfiability Modulo Theories (SMT) solver with limited success, and had to resort to culling the crossword grid based on candidate generations and ground truth answers.
In this paper, we undertake the challenge of solving a full crossword puzzle by developing an algorithm that incorporates grid information based on the answers given by an LLM. We find that with the application of our algorithm, the base performance of LLMs can be boosted significantly, and that the gap between the performance of LLMs and human experts is much smaller than previously anticipated.
@@ 3.1 Datasets
Most of our analysis is performed on three crossword puzzle datasets. The first one covering straight or American style crosswords was curated by us from the very popular and long running puzzle section of the NYT. The other two, namely, Cryptonite by Efrat et al. (2021) and word-init-disjoint by Rozner et al. (2021) cover cryptic crossword puzzles. As discussed in Section 2, the two cryptic crossword datasets have methodological differences in how they partition training-test splits, however, this is not a very pertinent consideration for us, since we do not perform any training. A bulk of the results are reported on the NYT dataset we curated for straight crosswords and word-init-disjoint (abbreviated as Init) for cryptic crosswords since it was found to be more challenging than Cryptonite for LLMs. See Appendix B for more details.
@@ 3.2 Answering Crossword Clues
The first step in solving a crossword puzzle is deciphering its clues, so following in the scheme of Rozner et al. (2021); Sadallah et al. (2024); Kulshreshtha et al. (2022) and Efrat et al. (2021) we first explore this as a QA task, where the clue and its length is provided to the LLM as a query, with the expectation that it will produce the corresponding answer.
@@
We find that the performance difference between 5-shot and 10-shot  [^3] We elected not to perform experiments with 10-shot prompts on Claude and GPT-4-Turbo due to budget constraints. prompted answers is not significant (see Figure 2), which is consistent with the findings by Sadallah et al. (2024). We also note that there is an appreciable difference (up to 5%) in performance of models across the board between Cryptonite and Init, with LLMs showing diminished performance on the Init dataset.
@@ 4 SweepClip - Our proposed algorithm to solve crosswords with LLMs
Our proposed algorithm (outlined in Algorithm 1 and detailed in Appendix A) first generates a set of candidate answers for all clues provided with the crossword (sweep) and uses a graph based criteria to eliminate answers that do not fit well to the grid (clip). Following this, we use the constraints generated from the previous step answers, generate more candidate answers, and prune the bad fitting candidates. We iteratively apply this strategy, until either - (i) the entire crossword is filled, (ii) we exceed a preset number of iterations, or (iii) we run out of LLM computational budget.
For pruning we use the largest connected component from the answers generated so far to ensure that the generated answers are somewhat coherent amongst themselves. This is over-restrictive, as it is possible that isolated answers are correct, however, we find that this strict pruning strategy works better to eliminate bad answers early rather than using the constraints imposed by potentially bad answers to generate further bad answers.
We would like to highlight that our algorithm is sub-optimal, and does not provide convergence or correctness guarantees beyond elimination of conflicting answers. Note that this problem is NP-Hard and it is difficult to find approximate solutions for it (Kulshreshtha et al., 2022). This algorithm was designed from the ground up to minimize cost incurred from calls to the LLM by avoiding branching based on LLM answers.
@@ 5.2 Solving Crossword Puzzles
For this task we employed our algorithm SweepClip on a set of 100 randomly sampled Monday NYT crossword puzzles. We used two LLMs for this - GPT-4-Turbo with max_depth of 30 and a budget of 0.5 USD per crossword, and LLaMA 3 70B with max_depth of 35 and a budget of 600 LLM calls.
@@
We find (see Table 4) that our algorithm with GPT-4-Turbo solves 48% of crosswords without any errors, and 55% of crosswords with at most 1 character wrong. The average character level accuracy in crossword solving is 93.1% ($\pm$ 14.1%). Our algorithm improves the clue-wise answer accuracy 888clue level accuracy is different from character level accuracy (Kulshreshtha et al., 2022), e.g. it is possible to have a filled in crossword without deciphering all clues. (exact match) to 89.6% ($\pm$ 16.9%) from the base accuracy (without the algorithm) which is 43.5% ($\pm$ 23.5%) an improvement of 2.1$\times$. Note the previously reported SoTA accuracy on this task was 26% with retrieval augmented generation and an SMT solver coupled with an oracle that eliminates parts of the crossword grid that do not have suitable generated answers (Kulshreshtha et al., 2022).
@@
Our algorithm implicitly performs self-consistency checks to improve candidate answers. For example the generated clue at the first sweep may be discarded and regenerated with a different prompt and then accepted if it does not produce conflicts, and is coherent with a larger number of other answers. Thus with the application of our algorithm, we have successfully exploited the constraint information to boost the performance of LLMs far beyond what would have been possible with straightforward QA like clue deciphering. To the best of our knowledge this is the first algorithm that demonstrates succesfully solving crosswords with the aid of LLMs.
@@ 5.3 Data Contamination and Generalizability
Having seen remarkable performance improvements in the current generation of SoTA LLMs, we set out to investigate the reason behind this success. To rule out data-contamination, we curated additional cryptic crossword clue datasets comprising entirely of puzzles published after May 20, 2024, which is after the knowledge cutoff date of all LLMs examined. These datasets are sourced from The Guardian [^9] https://www.theguardian.com/ and Lovatts Puzzles [^10] https://lovattspuzzles.com. Note, the Init dataset is also sourced from the The Guardian, thus the results reported in Table 5 should be consistent.  [^11] More details on data curation is in Appendix B.
@@
We see no appreciable difference in performance on the post-cutoff dataset, (see Table 5) leading us to conclude that these LLMs can generalize beyond potential contamination in their training set. The performance of models on the Lovatts Puzzles dataset (post-cutoff) is even better than that on the Guardian dataset.
To ascertain if the models can reason about unseen cryptic crossword clues to arrive at correct responses, we elect to perform human evaluation121212 Further details in Appendix E of model responses. We employ a 3-shot Chain-of-thought prompt to elicit a reasoned response to crossword clues from GPT-4-Turbo, which are then analyzed by our team for soundness vis-à-vis factual and logical errors. We choose 100 samples from the post-cutoff Lovatts set and rely on consensus of all evaluators to ascertain if a response contains factual or logical errors (e.g. wrong counting, wrong anagram, drawing a conclusion that does not follow from the premise).
@@ 6 Conclusions and Discussions
Solving crosswords requires proficiency in a multitude of desirable skills, and cryptic crosswords were previously thought to be firmly in the domain of human experts. Our findings challenge this notion. We have demonstrated that the current generation of SoTA LLMs show significantly improved aptitude at solving straight and cryptic crossword puzzle clues, and can exploit constraints provided by partially solved crosswords to boost this performance further. We also found that this emergent ability generalizes well to the post knowledge cut-off regime and is accompanied with the capacity to produce reasoned explanations.
@@
This calls into question previously reported results which claim that LLMs massively lag behind human experts at solving cryptic crosswords. We are firmly of the opinion that the current gap between human experts and LLMs will cease to exist in the near future. Incorporating length constraints effectively still remains a challenge, and further enhancements in numerical reasoning abilities in LLMs is a promising avenue to pursue.
@@
Limitation 2 - Algorithm Our algorithm is potentially sub-optimal as it discards potential correct answers and does not fully explore the consequences of each possible generated answer. This is again a computational cost saving technique. Ideally, we would generate one candidate answer, followed by all its neighbors, and whenever there is a conflict, branch out and explore all options to see which is a better fit. We elect not to do this, because it involves a potentially exponential number of LLM calls. Note that this problem is NP-Hard, however, future studies with much larger computational budget for LLM calls will definitely have more success employing a more thorough search strategy.
@@ C.1 General Overview
Phi 3 mini Instruct 3.8B 4K Oct. 2023
Mistral v0.2 Instruct 7B 32K Dec. 2023
LLaMA 3 Instruct 8B 8K Mar. 2023
Mixtral v0.1 8x7B 32K Dec. 2023
LLaMA 2 70B 4K Sep. 2022
LLaMA 3 70B 8K Dec. 2023
Claude 3 sonnet 20140229 ? 200K Mar. 2024
GPT-3.5-Turbo-0125 ? 16K Sep. 2021
@@
We used temperature=0.5 for all models and experiments. For few-shot prompt responses max-tokens were set to 10, and for Chain-of-thought responses max-tokens was set to 1000. top_p, top_k or any other parameters were not changed. The Jinja templates used to format the text prompts are provided alongside the paper, and are borrowed from the model-cards of the respective models. Claude and GPT models were accessed through their respective APIs and all other models were run locally on one server consisting of 2x80GB Nvidia A100 GPUs. All models were used in bf16 format whenever supported. Format of prompts are given in Appendix D.